How long would it take to sail 1,000 miles?
Found 2 solutions by oberobic, Maths68:
Answer by oberobic(2304)

It depends on the wind and the design of the sailboat.
If the sailboat can maintain 10 mph, then it would take 100 hrs.
If the sailboat can maintain 20 mph, then it would take 50 hrs.
But it depends...

Answer by Maths68(1474)

Its depend on your speed
If you are traveling 50 miles per hour (50mph) then = 1000/50=20 hours
or
If you are traveling 100 miles per hour (100mph) then = 1000/100=10 hours
Calculating such type of questions we usual use following formula
Distance= Speed * time
Above formula can be rearrange according to your requirment let say if you need to know the time
above formula will be
time=Distance/Speed
or
if you want to know Speed
Speed = Distance/time